Working alone, Ryan can dig a 10 ft by 10 ft
hole in five hours. Castel can dig the same
hole in six hours. How long would it take
them if they worked together?

I tried several ways to solve it but I can't seem to get the right answer. Because the answer to the problem is not what I am getting. Please I need help?

Answer by Alan3354(30983)

It's a matter of perspective.
Ryan digs the hole in 5 hours, which is 1/5 hole per hour.
Castel digs it in 6 hours, which is 1/6 hole per hour.
Together, they do the sum, 1/5 + 1/6 per hour = 11/30 holes per hour.
11/30 hole/hour --> 30/11 hours/hole
----------------
Use the shortcut, product/sum
5*6/(5+6) = 30/11